OpenStudy (kbonello):

Verify the trigonometric identity: (1 + tan^2x) cos^2x = 1

OpenStudy (vishweshshrimali5):

Okay... remember the identity... \(1 + \tan ^2 x = \)____

OpenStudy (vishweshshrimali5):

Can you tell me what comes in the blank?

OpenStudy (daniel.ohearn1):

It doesn't look right to me

OpenStudy (vishweshshrimali5):

See.. we have... \(1 + \tan^2 x = \sec^2 x\) And what is \(\sec x\) in terms of \(\cos x\) ?

OpenStudy (daniel.ohearn1):

cos(2x) = cos^2(x) -sin^2(x)

OpenStudy (vishweshshrimali5):

Looks like OP is offline... anyways I will give the final hint... \(\sec x = \cfrac{1}{\cos x}\) Using this and the identity I mentioned earlier you can get the answer

OpenStudy (faiqraees):

@daniel.ohearn1 It is cosine squared x

jhonyy9 (jhonyy9):

(1+tan^2 x)cos^2 x = 1 (1+sin^2 x/cos^2 x )cos^2 x = 1 ((cos^2 x +sin^2 x)/cos^2 x)cos^2 x = 1 (1/cos^2 x)cos^2 x = 1 1=1